OpenStudy (anonymous):

What are the steps to find the square root of - 121/196? The negative sign is confusing me.

OpenStudy (mimi_x3):

\[\sqrt{-\frac{121}{196}}\]?

OpenStudy (anonymous):

the negative sign was outside the square root sign, to the left of it.

OpenStudy (mimi_x3):

\[-\sqrt{\frac{121}{196}}\]?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

Oh. just ignore it for the moment.

OpenStudy (mimi_x3):

Just take the squareroot of \(\frac{121}{196}\) that you would normally do. Then add the minus sign at the end.

OpenStudy (mimi_x3):

Well, \(\sqrt{121} = 11\); \(\sqrt{196} = 14\)

OpenStudy (mimi_x3):

So, \[-\sqrt{\frac{121}{196}}= -\frac{11}{14}\]

OpenStudy (anonymous):

ah alright, thanks!

OpenStudy (mimi_x3):

:)